raw the combined series-parallel circuit as a simple circuit. Determine & include on Simple Circuit Schematic: Vi. I. Ry&P A series-parallel combination circuit (100 )(2SO)--11.43A oor 2S0 250 ? 350 ? R3 R, 200 ? 12 3 sor 2oo
Answer #1

Solution :-

1) Rp1=(R1R2/(R1+R2)~71.43 Ohms

Rp2=(R3R4/(R3+R4)~127.3 Ohms

Re=198.7 Ohms

2) I=U/Re=0.12 A

I=I1+I2=I3+I4

I1R1=I2R2

I3R3=I4R4

Solve this system of 3 eqs and get:

I2~0.035A

I4~0.086A

I3~0.044A

I4~0.076A

3) P=I2R

Replace the corresponding I in this formula and make the numerical calculation.

4) Sum of Pi (calculated at point 3) should get Ptot=I2Re~2.861 W